# How Numerical Precision Affects Mathematical Reasoning Capabilities of LLMs

### v1

@@ Abstract
Despite the remarkable success of Transformer-based Large Language Models (LLMs) across various domains, understanding and enhancing their mathematical capabilities remains a significant challenge. In this paper, we conduct a rigorous theoretical analysis of LLMs’ mathematical abilities, with a specific focus on their arithmetic performances. We identify numerical precision as a key factor that influences their effectiveness in mathematical tasks. Our results show that Transformers operating with low numerical precision fail to address arithmetic tasks, such as iterated addition and integer multiplication, unless the model size grows super-polynomially with respect to the input length. In contrast, Transformers with standard numerical precision can efficiently handle these tasks with significantly smaller model sizes. We further support our theoretical findings through empirical experiments that explore the impact of varying numerical precision on arithmetic tasks, providing valuable insights for improving the mathematical reasoning capabilities of LLMs.
How Numerical Precision Affects Mathematical Reasoning Capabilities of LLMs
Guhao Fengthanks: Equal contribution.1, Kai Yangfootnotemark: 1, Yuntian Gu1, Xinyue Ai1, Shengjie Luo1, Jiacheng Sun2, Di He1, Zhenguo Li2, Liwei Wang1 1Peking University   2Huawei Noah’s Ark Lab
@@ 1 Introduction
Integer Addition $\operatorname{ADD}(n,p)$ Constant $O(n^{2})$
Iterated Addition $\operatorname{IterADD}(n,k,p)$ Constant Super-polynomial
Integer Multiplication $\operatorname{Mul}(n,l,p)$ $O(n^{2})$ Super-polynomial
@@
Table 1: The model size w.r.t. the input size required for various arithmetic tasks on bounded-depth Transformers, under both standard and low numerical precision. Blue denotes the small model size, and red represents the large model size.
@@
Transformer-based Large Language Models (LLMs), such as GPT OpenAI (2023), Claude Anthropic (2024), and LLaMA Dubey et al. (2024), have achieved impressive performance across a broad range of natural language tasks Zhu et al. (2024); Basyal and Sanghvi (2023); Shao et al. (2023). Despite the great success, significant challenges remain when applying LLMs to mathematical problem-solving. Unlike many typical NLP tasks, which often depend on pattern recognition and statistical correlations Blei et al. (2003), mathematical reasoning requires rigorous logical deduction in a specific order (Frieder et al., 2024; Bubeck et al., 2023). To address these challenges, various strategies have been proposed, including carefully designed prompting strategies (Wei et al., 2022b; Yamauchi et al., 2023; Imani et al., 2023) and inference-based searching method (Kang et al., 2024; Wu et al., 2024a; Snell et al., 2024; Brown et al., 2024). However, a comprehensive understanding of the intrinsic limitations that restrict the mathematical reasoning capabilities of LLMs remains elusive.
@@
In principle, mathematical reasoning, built on basic arithmetical operations, requires the accurate computation of intermediate results throughout the reasoning process (Bubeck et al., 2023; Lee et al., 2024). There exist works (Feng et al., 2023; Yang et al., 2024) exploring the arithmetic capabilities of LLMs with Chain of Thought prompting (Wei et al., 2022b). However, unlike the tokenization methods of modern LLMs (Dubey et al., 2024; OpenAI, 2023), where each digit in a number is tokenized individually, these works treat a whole number as one token. Under this assumption, each distinct number occupies a unique position in the vocabulary, leading to an essential mismatch with practical implementations. Moreover, recent studies have demonstrated that LLMs operating with reduced numerical precision (e.g., int4) exhibit a significant decline in performance on mathematical tasks (Jin et al., 2024; Marchisio et al., 2024).
@@
In this paper, we provide a rigorous theoretical investigation of the arithmetical abilities of LLMs under the autoregressive paradigm. Specifically, we follow the tokenization method of modern LLMs, allowing the models to process and generate numbers digit by digit. Under these assumptions, we identify numerical precision as a key factor influencing their performance in arithmetical tasks. Our analysis focuses on three elementary arithmetic tasks: integer addition, iterated addition, and integer multiplication, which serve as elementary building blocks in solving complex real-world math problems.
@@
To elucidate the role of numerical precision, we first examine the expressiveness of Transformers operating under low precision, such as int8 and int4. We establish foundational impossibility results for low-precision Transformers, demonstrating that such models require super-polynomial size with respect to input length to solve iterated addition and integer multiplication (Theorems 4.2 and 4.3). Our proofs, grounded in complexity theory (Arora and Barak, 2009; Razborov, 1987), show that this limitation arises from the inability of individual neurons to store intermediate results during arithmetic computations. As a result, a significantly larger number of neurons is required to distribute the computation and avoid overflow.
@@
We further demonstrate that increasing numerical precision is essential to overcoming this limitation: as the precision increases, the model size required to solve these arithmetic problems decreases dramatically. In particular, we prove that a bounded-depth Transformer with standard precision, such as float32, can solve all three tasks efficiently, requiring constant model size for integer and iterated addition (Theorems 5.1 and 5.2), and quadric size for integer multiplication (Theorem 5.3). At standard precision, it is sufficient for LLMs to accurately store relatively large numbers within individual neurons, enabling the Transformer to perform arithmetic tasks more effectively. These results not only underscore the critical role of numerical precision in enhancing the theoretical capabilities of LLMs for mathematical tasks but also offer practical insights. While low-precision models may provide computational efficiency, ensuring adequate numerical precision is essential for applications involving complex mathematical reasoning. A summary of our results is presented in Table 1.
@@
In addition to theoretical analysis, we conduct extensive experiments to empirically validate our conclusions. First, we evaluate the performance of Transformers trained from scratch on the aforementioned arithmetic tasks, systematically examining how problem size and numerical precision impact their capabilities. Furthermore, we also conduct experiments on LLAMA 3.1 (Dubey et al., 2024) to evaluate the performance of these arithmetic tasks under different numerical precision. Our empirical results demonstrate that both low-precision and standard-precision Transformers perform adequately on the integer addition task. However, as task complexity increases—particularly in iterated addition and integer multiplication—the decrease in precision in Transformers results in significant performance degradation. These findings align with our theoretical predictions and offer practical guidance for enhancing LLM performance in mathematical reasoning tasks.
@@ 2 Preliminary
where ${\bm{W}}_{\text{Q}}^{(l,h)},{\bm{W}}_{\text{K}}^{(l,h)},{\bm{W}}_{\text{V}}^{(l,h)},{\bm{W}}_{\text{O}}^{(l,h)}\in\mathbb{R}^{\lceil\frac{d}{H}\rceil\times d}$ are the query, key, value, and output matrices of the $h$-th head in the $l$-th layer. The weight matrices in the feed-forward network are denoted as ${\bm{W}}_{1}^{(l)},{\bm{W}}_{2}^{(l)}\in\mathbb{R}^{d\times d}$. The activation function $\sigma$ is chosen to be GeLU (Hendrycks and Gimpel, 2016), following the work of Radford et al. (2019); Devlin et al. (2019).
@@ 3 Problem Setup
Integer Representations.
We assume all integers are non-negative and represented in the base-$p$ form, where $p\geq 2$ is a fixed base. Specifically, an integer with $n$ digits is expressed as $(x_{n-1}\cdots x_{0})_{p}$, where the sequence of digits $\mathbf{x}=[x_{n-1},\cdots,x_{0}]$ is the input to the Transformer model. Each digit $x_{i}$ is treated as an individual token according to the model’s tokenization scheme. During generation, the Transformer also produces the outcome digit by digit, respecting the established tokenization structure.
Unlike prior works (Feng et al., 2023; Yang et al., 2024), which represent entire integers as single tokens, we treat each digit of an integer as an individual token, allowing Transformers to process and generate numbers digit by digit. Given that most modern LLMs tokenize large numbers in this manner, our approach not only aligns with practical tokenization strategies but also mitigates the issue of inflated vocabulary size.
Integer Addition.
Let ${\bm{a}}=(a_{n_{1}-1}\cdots a_{0})_{p}$ and ${\bm{b}}=(b_{n_{2}-1}\cdots b_{0})_{p}$ denote two integers encoded in base-$p$. The input sequence is constructed by concatenating the tokenized representations of ${\bm{a}}$ and ${\bm{b}}$, separated by the addition operator ‘$+$’. The task is to generate the sum ${\bm{s}}=(s_{n}\cdots s_{0})_{p}$, which contains $n+1$ digits where $n=\max(n_{1},n_{2})$ and represents the base-$p$ sum of ${\bm{a}}$ and ${\bm{b}}$, output digit-by-digit.
Iterated Addition.
Consider $k$ integers ${\bm{a}}_{1}=(a_{1,n_{1}-1}\cdots a_{1,0})_{p},\;\cdots,\;{\bm{a}}_{k}=(a_{k,n_{k}-1}\cdots a_{k,0})_{p}$ where $n=\max\{n_{1},\cdots,n_{k}\}$. The input sequence consists of the tokenized representations of these integers, concatenated and separated by the addition operator ‘$+$’. The Transformer model must output the sum ${\bm{s}}=(s_{n+\lceil\log_{p}k\rceil-1}\cdots s_{0})_{p}$, which contains $n+\lceil\log_{p}k\rceil$ digits and represents the base-$p$ sum of the $k$ integers.
Integer Multiplication.
@@
In the integer multiplication task, the goal is to compute the product of two integers, truncated to a pre-specified length $l$. Let ${\bm{a}}=(a_{n_{1}-1}\cdots a_{0})_{p}$ and ${\bm{b}}=(b_{n_{2}-1}\cdots b_{0})_{p}$ be two integers, each no longer than $n$ digits in base-$p$. The input sequence is constructed by concatenating the tokenized representations of ${\bm{a}}$ and ${\bm{b}}$, separated by the multiplication operator ‘$\times$’. The objective is to generate the lowest $l$ digits ($l\leq 2n$) of the product ${\bm{s}}=(s_{l-1}\cdots s_{0})_{p}$ in base-$p$ arithmetic.
@@
We consider a generalized case of integer multiplication where overflow may occur if the result exceeds the given digit length. Standard integer multiplication is a special case of this framework when $l=2n$.
Figure 1 illustrates examples of the three tasks. Integer addition is the simplest of these operations and can be viewed as a specific instance of iterated addition. Furthermore, integer multiplication inherently involves the summation of several intermediate products. Consequently, we present these tasks in increasing order of complexity. In the subsequent sections, we use the notations $\operatorname{ADD}(n,p)$ to denote $n$-digit addition in base-$p$ arithmetic, $\operatorname{IterADD}(n,k,p)$ for the iterated addition of $k$ integers with $n$ digits each in base-$p$, and $\operatorname{MUL}(n,l,p)$ for the multiplication of two $n$-digit integers in base-$p$, truncated to $l$ digits.
@@ 4 Low-Precision Transformers Struggle with Basic Arithmetic Tasks
To formalize these limitations, we build on the framework introduced by Li et al. (2024) and utilize the setting of a constant-precision Transformer (See formal definition in Appendix B). In this setting, the internal states of the model’s neurons are constrained to represent real numbers using only $c$ bits, where $c$ is a small constant independent of the input sequence length. These numbers may be represented by floating point in IEEE 754 formats (Kahan, 1996) or fixed point formats. This configuration mirrors many practical deployment scenarios, in which LLMs often employ reduced-precision formats such as float8, int8, or even int4, particularly during inference Han et al. (2015). Given that these models typically process input sequences comprising thousands of tokens, it is reasonable and realistic to assume that the numerical precision remains fixed at a small constant, independent of sequence length. Under the constant-precision setting, we examine the expressiveness of the Transformer model in elementary arithmetic problems.
@@
For any fixed integers $p$, there exist constant-precision Transformers with constant depth $L$ and hidden dimension $d=O(n^{2})$ that can solve the $\operatorname{ADD}(n,p)$ task.
Theorem 4.1 suggests that the bounded-depth Transformers with reasonable hidden dimensions are capable of solving the integer addition task. However, as we will show in subsequent theorems, constant-precision Transformers exhibit pronounced limitations when considering more complex arithmetic problems.
@@
For any integers $p$ and $L$, and for any polynomial $f$, there exist problem scales $n$ and $k$ such that no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d<f(n,k)$ can correctly solve the $\operatorname{IterADD}(n,p)$ task.
@@
For any integers $p$ and $L$, and for any polynomial $f$, there exist problem scales $n$ and $l$ such that no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d<f(n,l)$ can correctly solve the $\operatorname{MUL}(n,l,p)$ task.
What accounts for this limitation? As presented in Appendix D, our proof is grounded in circuit complexity theory. By modeling the constant-precision Transformer as a computational circuit, we rigorously analyze its expressive limitations through the lens of circuit complexity (Feng et al., 2023; Merrill et al., 2022; Merrill and Sabharwal, 2023; Li et al., 2024). Specifically, Li et al. (2024) proves that the expressiveness of constant-precision Transformers with polynomial size and bounded depth is upper-bounded by the computation complexity class $\mathsf{AC}^{0}$. In contrast, we demonstrate that the complexity of tasks such as $\operatorname{IterADD}$ and $\operatorname{MUL}$ exceeds that of $\mathsf{AC}^{0}$, using reductions from Majority, a well-established problem that has been provably unsolvable by the circuits in $\mathsf{AC}^{0}$ (Razborov, 1987; Smolensky, 1987). Consequently, these tasks are inherently hard for low-precision Transformers.
Practical Implications. Intuitively, when performing iterated addition, a Transformer must compute the sum of multiple digits at one time. However, under low-precision constraints, maintaining accurate intermediate results becomes significantly challenging. For example, for the model under int4 precision, adding $8$ and $9$ in base-$10$ arithmetic will require several neurons to store the results to avoid overflow. In such cases, a Transformer operating in a low-precision regime would either require additional steps to compute the correct sum or need much more neurons to distribute the computation to avoid overflow. This will finally result in the Transformers failing to complete the task with a fixed number of layers and a reasonable model size.
While low-precision Transformers can effectively handle some simplest arithmetic tasks, such as basic integer addition, their capacity is severely limited when addressing more complex tasks. As demonstrated, low numerical precision, such as int8 and float8, imposes fundamental constraints, preventing these models from solving problems that would require Transformers with super-polynomial size.

### v2

# How Numerical Precision Affects Arithmetical Reasoning Capabilities of LLMs

@@ Abstract
Despite the remarkable success of Transformer-based large language models (LLMs) across various domains, understanding and enhancing their mathematical capabilities remains a significant challenge. In this paper, we conduct a rigorous theoretical analysis of LLMs’ mathematical abilities, with a specific focus on their arithmetic performances. We identify numerical precision as a key factor that influences their effectiveness in arithmetical tasks. Our results show that Transformers operating with low numerical precision fail to address arithmetic tasks, such as iterated addition and integer multiplication, unless the model size grows super-polynomially with respect to the input length. In contrast, Transformers with standard numerical precision can efficiently handle these tasks with significantly smaller model sizes. We further support our theoretical findings through empirical experiments that explore the impact of varying numerical precision on arithmetic tasks, providing valuable insights for improving the mathematical reasoning capabilities of LLMs.
How Numerical Precision Affects Arithmetical Reasoning Capabilities of LLMs
Guhao Fengthanks: Equal contribution.1, Kai Yangfootnotemark: 2, Yuntian Gu1, Xinyue Ai2, Shengjie Luo1, Jiacheng Sun3, Di He1, Zhenguo Li3, Liwei Wang1,4 1State Key Laboratory of General Artificial Intelligence, School of Intelligence Science and Technology, Peking University 2School of EECS, Peking University   3Huawei Noah’s Ark Lab 4Center for Machine Learning Research, Peking University Correspondence: dihe@pku.edu.cn
@@ 1 Introduction
Integer Addition $\operatorname{ADD}_{p}(n)$ Constant $O(n^{2})$
Iterated Addition $\operatorname{IterADD}_{p}(n,k)$ Constant Super-polynomial
Integer Multiplication $\operatorname{Mul}_{p}(n,l)$ $O(n^{2})$ Super-polynomial
@@
Table 1: The model size w.r.t. the input size required for various arithmetic tasks on bounded-depth Transformers, under both standard and low numerical precision. Blue denotes the acceptable model size, and red represents the unaffordable model size.
@@
Transformer-based LLMs, such as GPT OpenAI (2023), Claude Anthropic (2024), and LLAMA Dubey et al. (2024), have achieved impressive performance across a broad range of natural language tasks Basyal and Sanghvi (2023); Shao et al. (2023); Zhu et al. (2024). Despite the great success, significant challenges remain when applying LLMs to mathematical problem-solving. Unlike many typical NLP tasks, which often depend on pattern recognition and statistical correlations Blei et al. (2003), mathematical reasoning requires rigorous logical deduction in a specific order (Bubeck et al., 2023; Frieder et al., 2024). To address these challenges, various strategies have been proposed, including carefully designed prompting strategies (Wei et al., 2022b; Yamauchi et al., 2023; Imani et al., 2023) and inference-based searching method (Kang et al., 2024; Wu et al., 2024a; Snell et al., 2024; Brown et al., 2024). However, a comprehensive understanding of the intrinsic limitations that restrict the mathematical reasoning capabilities of LLMs remains elusive.
@@
In principle, mathematical reasoning, built on basic arithmetical operations, requires accurate computation of intermediate results throughout the reasoning process (Bubeck et al., 2023; Lee et al., 2024). There exist works (Feng et al., 2023; Yang et al., 2024) exploring the arithmetic capabilities of LLMs with Chain of Thought (CoT) prompting (Wei et al., 2022b). However, these investigations often deviate from the tokenization strategies employed by modern LLMs (OpenAI, 2023; Dubey et al., 2024), where numbers are typically segmented into tokens of at most three digits. Under the assumption of Feng et al. (2023) and Yang et al. (2024), each distinct number occupies a unique position in the vocabulary, leading to an essential mismatch with practical implementations. Moreover, recent studies have demonstrated that LLMs operating with reduced numerical precision (e.g., int4) exhibit a significant decline in performance on mathematical tasks (Jin et al., 2024; Marchisio et al., 2024).
@@
In this paper, we provide a rigorous theoretical investigation of the arithmetical abilities of LLMs under the autoregressive paradigm. Specifically, we adopt the tokenization approach used in modern LLMs, where numbers are processed and generated token by token, with each token representing only a small number of digits. Under these assumptions, we identify numerical precision as a key factor influencing their performance in arithmetical tasks. Our analysis focuses on three elementary arithmetic tasks: integer addition, iterated addition, and integer multiplication, which serve as elementary building blocks in solving complex real-world math problems.
@@
To elucidate the role of numerical precision, we first examine the expressiveness of Transformers operating under low precision, such as int8 and int4. We establish foundational impossibility results for low-precision Transformers, demonstrating that such models require super-polynomial size with respect to input length to solve iterated addition and integer multiplication (Theorems 4.2 and 4.3). Our proofs, grounded in complexity theory (Razborov, 1987; Arora and Barak, 2009), show that this limitation arises from the inability of individual neurons to store intermediate results during arithmetic computations. As a result, a significantly larger number of neurons is required to distribute the computation and avoid overflow.
@@
We further demonstrate that increasing numerical precision is essential to addressing this limitation. Specifically, as numerical precision improves, the model size required to solve arithmetic tasks decreases significantly. In particular, we prove that a bounded-depth Transformer operating with standard precision (e.g., float32) can efficiently and reliably solve all three tasks under consideration. For both integer and iterated addition, the required model size remains constant and independent of the input length (Theorems 5.1 and 5.2), while for integer multiplication, the model size scales quadratically w.r.t the input length (Theorem 5.3). These results highlight that standard numerical precision is sufficient for LLMs to effectively perform arithmetic tasks. Our findings emphasize the practical importance of numerical precision in mathematical reasoning. While low-precision models may offer computational advantages, ensuring sufficient numerical precision is critical for tasks involving complex arithmetic. A summary of our main results is provided in Table 1.
@@
In addition to theoretical analysis, we conduct extensive experiments to validate our conclusions. First, we evaluate the performance of Transformers trained from scratch on the aforementioned arithmetic tasks, systematically examining how problem size and numerical precision impact their capabilities. Furthermore, we also conduct experiments on LLAMA-3.1-8B Instruct (Dubey et al., 2024) to evaluate the performance of these arithmetic tasks under different numerical precision. Our empirical results demonstrate that both low-precision and standard-precision Transformers perform adequately on the integer addition task. However, as task complexity increases—particularly in iterated addition and integer multiplication—the decrease in precision in Transformers results in significant performance degradation. These findings align with our theoretical predictions and offer practical guidance for enhancing LLM performance in mathematical reasoning tasks.
@@ 2 Preliminary
where ${\bm{W}}_{\text{Q}}^{(l,h)}$,${\bm{W}}_{\text{K}}^{(l,h)}$,${\bm{W}}_{\text{V}}^{(l,h)}$,${\bm{W}}_{\text{O}}^{(l,h)}$ $\in\mathbb{R}^{\lceil\frac{d}{H}\rceil\times d}$ are the query, key, value, and output matrices of the $h$-th head in the $l$-th layer. The weight matrices in the feed-forward network are denoted as ${\bm{W}}_{1}^{(l)},{\bm{W}}_{2}^{(l)}\in\mathbb{R}^{d\times d}$. The activation function $\sigma$ is chosen to be GeLU (Hendrycks and Gimpel, 2016), following the work of Radford et al. (2019).
@@ 3 Problem Setup
Integer Representation and Tokenization. We consider all integers to be non-negative and represented in base-$p$ notation, where $p\geq 2$ is a fixed base. Specifically, an integer with $n$ digits is expressed as $(x_{n-1}\cdots x_{0})_{p}$. To tokenize this sequence, we employ a tokenizer, denoted by ${\bm{T}}_{c}$, that partitions $\mathbf{x}$ into tokens, each containing at most $c$ contiguous digits. Formally, let the sequence ${\bm{t}}=[t_{k-1},\ldots,t_{0}]={\bm{T}}_{c}([x_{n-1},\ldots,x_{0}])$, where $k=\lceil\frac{n}{c}\rceil$ we have
$t_{i}=\begin{cases}[x_{ic},x_{ic+1},\cdots,x_{ic+c-1}],&i<k-1;\\ [x_{ic},x_{ic+1},\cdots,x_{n-1}],&i=k-1.\\ \end{cases}$ (1)
During sequence generation, the Transformer model outputs the target tokens sequentially, strictly following the tokenization scheme of tokenizer ${\bm{T}}_{c}$. Unlike prior works (Feng et al., 2023; Yang et al., 2024), which represent entire integers as single tokens, this approach aligns with prevalent tokenization strategies employed by modern LLMs (OpenAI, 2023; Dubey et al., 2024) and enables Transformers to process and generate numbers token by token. Further discussion and illustrative examples of the tokenization scheme are provided in Section B.4.
Integer Addition. Let ${\bm{a}}=(a_{n_{1}-1}\cdots a_{0})_{p}$ and ${\bm{b}}=(b_{n_{2}-1}\cdots b_{0})_{p}$ denote two integers represented in base-$p$. Their sum is expressed as ${\bm{s}}=(s_{n}\cdots s_{0})_{p}={\bm{a}}+{\bm{b}}$. Let ${\bm{T}}_{c}$ represent the tokenizer. The input sequence is constructed by concatenating the tokenized representations of ${\bm{a}}$ and ${\bm{b}}$, i.e., ${\bm{T}}_{c}({\bm{a}})$ and ${\bm{T}}_{c}({\bm{b}})$, with the addition operator token ‘$+$’ placed between them, and the equality operator token ‘$=$’ appended at the end. The task is to generate the tokenized representation of the result, ${\bm{T}}_{c}({\bm{s}})$, sequentially, one token at a time.
Iterated Addition. Now consider $k$ integers in base-$p$: ${\bm{a}}_{1}=(a_{1,n_{1}-1}\cdots a_{1,0})_{p},\;\ldots,\;{\bm{a}}_{k}=(a_{k,n_{k}-1}\cdots a_{k,0})_{p}$, where $n=\max\{n_{1},\ldots,n_{k}\}$. Their sum is denoted as ${\bm{s}}=(s_{n-1}\cdots s_{0})_{p}=\sum_{i\in[k]}{\bm{a}}_{i}$, where $n=\max_{i\in[k]}\{n_{k}\}+\lceil\log k\rceil$. Let ${\bm{T}}_{c}$ denote the tokenizer. The input sequence is formed by concatenating the tokenized representations of these integers, separated by the addition operator token ‘$+$’, followed by the equality operator token ‘$=$’ appended at the end. The objective is for the Transformer to generate the tokenized representation of the sum, ${\bm{T}}_{c}({\bm{s}})$, sequentially, one token at a time.
@@
Integer Multiplication. The integer multiplication task involves computing the product of two integers, truncated to a predefined length $l$. Let ${\bm{a}}=(a_{n_{1}-1}\cdots a_{0})_{p}$ and ${\bm{b}}=(b_{n_{2}-1}\cdots b_{0})_{p}$ represent two integers in base-$p$, and let $n=\max\{n_{1},n_{2}\}$. Their product is given by ${\bm{s}}=(s_{2n-1}\cdots s_{0})_{p}={\bm{a}}\times{\bm{b}}$. Let ${\bm{T}}_{c}$ denote the tokenizer. The input sequence is constructed by concatenating the tokenized representations of ${\bm{a}}$ and ${\bm{b}}$, separated by the multiplication operator token ‘$\times$’, with the equality operator token ‘$=$’ appended at the end. The objective is to generate the tokenized representation of the product’s least significant $l$ digits, ${\bm{T}}_{c}([s_{l-1},s_{l-2},\ldots,s_{0}])$, where $l\leq 2n$.
@@
We consider a generalized case of integer multiplication where overflow may occur if the result exceeds the given digit length. Standard integer multiplication is a special case of this framework when $l=n_{1}+n_{2}$.
Figure 1 presents examples of these tasks. Integer addition is the simplest of these tasks and can be viewed as a specific instance of iterated addition. Furthermore, integer multiplication inherently involves the summation of several intermediate products. Consequently, we present these tasks in increasing order of complexity. In the subsequent sections, we use the notations $\operatorname{ADD}_{p}(n)$ to denote addition with at most $n$ digits in base-$p$ arithmetic, $\operatorname{IterADD}_{p}(n,k)$ for the iterated addition of $k$ integers with at most $n$ digits each in base-$p$, and $\operatorname{MUL}_{p}(n,l)$ for the multiplication of two integers with at most $n$ digits in base-$p$, truncated to $l$ digits.
@@ 4 Low-Precision Transformers Struggle with Basic Arithmetic Tasks
To formalize these limitations, we build on the framework introduced by Li et al. (2024) and utilize the setting of a constant-precision Transformer (See formal definition in Section B.2). In this setting, the internal states of the model’s neurons are constrained to represent real numbers using only $c$ bits, where $c$ is a small constant independent of the input sequence length. These numbers may be represented by floating point in IEEE 754 formats (Kahan, 1996) or fixed point formats. This configuration mirrors many practical deployment scenarios, in which LLMs often employ reduced-precision formats such as float8, int8, or even int4, particularly during inference Han et al. (2015). Given that these models typically process input sequences comprising thousands of tokens, it is reasonable and realistic to assume that the numerical precision remains fixed at a small constant, independent of sequence length. Under the constant-precision setting, we examine the expressiveness of the Transformer model in elementary arithmetic problems.
@@
Fix integers $p\geq 2$ and $c\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. There exist constant-precision Transformers with constant depth (independent of $n$) and hidden dimension $d=O(n^{2})$ that can solve the $\operatorname{ADD}_{p}(n)$ task.
Theorem 4.1 suggests that the bounded-depth Transformers with reasonable hidden dimensions are capable of solving the integer addition task. However, as we will show in subsequent theorems, constant-precision Transformers exhibit pronounced limitations when considering more complex arithmetic problems. For the page limitation, we give the detailed proof of Theorem 4.1 in Section D.1.
@@
Fix integers $p\geq 2$ and $c,L\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. For any polynomial $f$, there exist problem scales $n$ and $k$ such that no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d<f(n,k)$ can correctly solve the $\operatorname{IterADD}_{p}(n,k)$ task.
@@
Fix integers $p\geq 2$ and $c,L\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. For any polynomial $f$, there exist problem scales $n$ and $l$ such that no constant-precision autoregressive Transformer with $L$ layers and hidden dimension $d<f(n,l)$ can correctly solve the $\operatorname{MUL}_{p}(n,l)$ task.
The detailed proof of Theorems 4.2 and 4.3 are presented in Sections D.2 and D.3.
What accounts for this limitation? As presented in Appendix D, our proof is grounded in circuit complexity theory. By modeling the constant-precision Transformer as a computational circuit, we rigorously analyze its expressive limitations through the lens of circuit complexity (Merrill et al., 2022; Merrill and Sabharwal, 2023; Feng et al., 2023; Li et al., 2024). Specifically, Li et al. (2024) proves that the expressiveness of constant-precision Transformers with polynomial size and bounded depth is upper-bounded by the computation complexity class $\mathsf{AC}^{0}$. In contrast, we demonstrate that the complexity of tasks such as $\operatorname{IterADD}$ and $\operatorname{MUL}$ exceeds that of $\mathsf{AC}^{0}$, using reductions from Majority, a well-established problem that has been provably unsolvable by the circuits in $\mathsf{AC}^{0}$ (Razborov, 1987; Smolensky, 1987). Consequently, these tasks are inherently hard for low-precision Transformers.
Practical Implications. While low-precision Transformers can effectively handle some of the simplest arithmetic tasks, such as basic integer addition, their capacity is severely limited when addressing more complex tasks. As demonstrated, low numerical precision, such as int4 and float8, imposes fundamental constraints, preventing these models from solving problems that would require Transformers with super-polynomial size.